# Long-range Meta-path Search on Large-scale Heterogeneous Graphs

### v4

@@ D.5 Further Comparison on Freebase Dataset
D.6 Ablation Study
Two components differentiate our LMSPS from other HGNNs: search algorithm and semantic fusion without attention. The search algorithm consists of a progressive sampling algorithm and a sampling evaluation strategy. Additionally, many HGNNs (Hu et al., 2020b; Yang et al., 2023; Mao et al., 2023) employ the attention mechanism to fuse semantic information. In LMSPS, the difference of importance between the searched effective meta-paths is much smaller than that between the full meta-path set, making the attention mechanism seem unnecessary. We explore how each of them improves performance through ablation studies on DBLP, IMDB, ACM, and ogbn-mag under maximum hops $6$, $5$, $5$, $6$, respectively. As shown in Table 11, the performance of LMSPS significantly decreases when removing progressive sampling or sampling evaluation strategy. Using a transformer block for semantic attention on searched meta-paths can not improve performance, indicating semantic attention is not necessary for LMSPS. In addition, employing a transformer block for semantic attention on all meta-paths shows slightly worse performance, even if using many more meta-paths, indicating that the transformer can’t eliminate the negative effects of noise. It is reasonable because attention values are calculated based on softmax and are positive even for noise. It can explain why transformer-based methods like HGT (Hu et al., 2020b) don’t work well on obgn-mag,
@@ D.6 Ablation Study
D.7 Hyperparameter Study
LMSPS randomly samples $M$ meta-paths at each epoch in the search stage and selects the top-$M$ meta-paths in the training stage. Here, we perform analysis on hyper-parameter $M$ on DBLP, IMDB, and ACM. As illustrated in Figure 5, the performance of LMSPS increases with the growth of $M$ when $M$ is small. In addition, the performance on DBLP and ACM slightly decreases when $M$ is larger than a certain threshold, indicating excessive meta-paths don’t benefit performance for some datasets. For unity, we set $M=30$ for all datasets.
@@ D.7 Hyperparameter Study
Appendix E Reproducibility Statement
We have provided the motivation for dataset selection and listed hyper-parameters used for the experiments. All reported results are the average of multiple experiments with standard deviations. We have included a pseudocode description of our method in Appendix B, and describe raining details in Section 6.2 and Appendix D.2. The source code has been submitted as supplementary materials with clear commands on reproducing our results.
Appendix F Limitations
It is noteworthy that the performance of LMSPS does not always increase with the value of the maximum hop. For instance, based on Figure 3 (a), LMSPS can effectively utilize 12-hop meta-paths on DBLP with high performance and low cost. However, the optimal performance for LMSPS is observed when the maximum hop is $6$. It can be attributed to the constraint of maintaining constant time complexity by setting the number of searched meta-paths to $30$ for different maximum hop values. When the maximum hop value is $12$, the number of target-node-related meta-paths exceeds $1400$ (Figure 3 (a)). Searching for $30$ effective meta-paths from such an extensive search space is notably challenging, despite LMSPS being the most effective method for meta-path search (as indicated in Table 9). The advantages of utilizing longer meta-path confers fall short of compensating for the associated drawbacks of much more challenging search space. So, the best maximum hop depends on the dataset and task and cannot be determined automatically.
Nevertheless, based on Table 1, the time complexity of LMSPS does not increase with the maximum hop. Consequently, it provides an effective solution for utilizing long-range dependency on heterogeneous graphs for possible applications on sparser real-world datasets. In Table 4, we conduct experiments on four constructed datasets based on ogbn-mag to demonstrate that the advantages of utilizing long-range dependencies are more obvious for sparser large-scale heterogeneous graphs. For the constructed datasets, we carefully avoid inappropriate preference seed settings of randomly removing by limiting the maximum in-degree related to edge type. We expect sparser large-scale real heterogeneous datasets or more suitable tasks requiring long-range dependency to emerge in the future so we can fully explore our method.

### v6

@@ E.3 Convergence of the Search Stage
E.4 Hyperparameter Study
LMSPS randomly samples $M$ meta-paths at each epoch in the search stage and selects the top-$M$ meta-paths in the training stage. Here, we perform analysis on hyper-parameter $M$ on DBLP, IMDB, and ACM. Freebase and OGBN-MAG are ignored because they are relatively large and the experiments are time-consuming. As illustrated in Figure 5, the performance of LMSPS increases with the growth of $M$ when $M$ is small. In addition, the performance on DBLP and ACM slightly decreases when $M$ is larger than a certain threshold, indicating excessive meta-paths don’t benefit performance for some datasets. For unity, we set $M=30$ for all datasets.
@@ E.4 Hyperparameter Study
Appendix F Reproducibility Statement
We have provided the details of datasets, tasks, baselines, and parameter settings in Appendix A and B and conducted the hyperparameter study in Appendix E.4. All reported results are the average of multiple experiments with standard deviations. We have included a pseudocode description of our method in Appendix C. The source code has been provided through an anonymized URL with clear commands on reproducing our results.
Appendix G Limitations
It is noteworthy that the performance of LMSPS does not always increase with the value of the maximum hop. For instance, based on Figure 3 (a), LMSPS can effectively utilize 12-hop meta-paths on DBLP with high performance and low cost. However, the optimal performance for LMSPS is observed when the maximum hop is $6$. It can be attributed to the constraint of maintaining constant time complexity by setting the number of searched meta-paths to $30$ for different maximum hop values. When the maximum hop value is $12$, the number of target-node-related meta-paths exceeds $1400$ (Figure 3 (a)). Searching for $30$ effective meta-paths from such an extensive search space is notably challenging, despite LMSPS being the most effective method for meta-path search (as indicated in Table 3). The benefits of using longer meta-paths are outweighed by the drawbacks of the significantly more complex search space. So, the best maximum hop depends on the dataset and task and cannot be determined automatically.
Nevertheless, based on Table 8, the time complexity of LMSPS does not increase with the maximum hop. Consequently, it provides an effective solution for utilizing long-range dependency on heterogeneous graphs for possible applications on sparser real-world datasets. In Table 5, we conduct experiments on four constructed datasets based on OGBN-MAG to demonstrate that the advantages of utilizing long-range dependencies are more obvious for sparser large-scale heterogeneous graphs. For the constructed datasets, we carefully avoid inappropriate preference seed settings of randomly removing by limiting the maximum in-degree related to edge type. We expect sparser large-scale real heterogeneous datasets or more suitable tasks requiring long-range dependency to emerge in the future so we can fully explore our approach.